You have $1,000,000 to invest in a stock portfolio. Your choices are Stock X with an expected return of 18 percent and Stock Y with an expected return of 10 percent. Your goal is to create a portfolio with an expected return of 13 percent. All money must be invested. How much will you invest in Stock X?
Answer #1

How much will you invest in Stock X?

$                 3,75,000
  • The Expected Return of a portfolio is the weighted average return of the individual securities in the portfolio, where the weights are the proportions invested in each security.
  • Let P be the proportion of Stock X in the portfolio. Then the proportion of Stock Y in the portfolio is 1-P
Stock Expected return (%) Proportion Expected return of portfolio (%)
a b c = a × b
X 18 P 18P
Y 10 1-P 10(1-P)
Expected return of portfolio (target) 13
  • 18P + 10(1-P) = 13
  • 18P + 10 - 10P = 13
  • 8P = 13 - 10
  • P = 3 ÷ 8 = 0.375
  • Therefore, proportion of stock X & Y:
    • X P 0.375 37.5%
      Y 1-P 1 - 0.375 = 0.625 62.5%
  • Verification:
    • Stock Expected return (%) Proportion Expected return of portfolio (%)
      a b c = a × b
      X 18 0.375 6.75
      Y 10 0.625 6.25
      Expected return of portfolio (target) 13.00
  • HOW MUCH TO INVEST IN EACH STOCK:
    • Stock Computation Amount invested
      X $1,000,000 × 37.5% $                 375,000
      Y $1,000,000 × 62.5% $                 625,000
      Total $               1,000,000